School district A has an assessed valuation of taxable property of $49,410,00. It has 5,400 public school students. School District B has an assessed valuation of taxable property of 86,260,000 and 9,500 public school pupils. which district has the greater ability to support its school? how much greater?
Answer #1

Solution: Total school going students = 5400 + 9500 = 14900

Proportion of school going students in district A = 5400 / 14900 = .3624

Proportion of school going students in district B = 9500 / 14900 = ..6375

Ability of district A = 4941000 * .3624 = 1790618.4

Ability of district B = 86,260,000 * .6375 = 54990750

So , district B has greater ability by (54990750 - 1790618.4 = 53200131.6)